OpenStudy (ilovebmth1234):

Consider this reaction: N2 (g) + O2 (g) <--> 2NO(g) What will happen if pressure is increased on the system? A) there will be a shift toward the products B) there will be a shift toward the reactants C) there will not be a shift D) Keq will increase

OpenStudy (ilovebmth1234):

@Awolflover1

OpenStudy (ilovebmth1234):

@just_one_last_goodbye

OpenStudy (n8kd_da_beast):

Wow I haven't seen these formulas in two years. I need to brush up on them.

NetflixAndNo (netflixandno):

Moles product = 2 Moles reactants = 1 + 1 = 2 Delta T = 2 - 2 = 0 There not be a shift.

NetflixAndNo (netflixandno):

There will not be a shift**

OpenStudy (ilovebmth1234):

so C?

NetflixAndNo (netflixandno):

Yes.

OpenStudy (ilovebmth1234):

thanks!! could you help me on like 2 more?

NetflixAndNo (netflixandno):

I can try, you're welcome c:

OpenStudy (ilovebmth1234):

Why is a chemical equilibrium described as dynamic? A) Because the reactant and product molecules are in constant motion. B) Because maximum randomness has been achieved. C) Because the pressure and temperature do not change. D) Because the concentrations of chemical species remain constant.

NetflixAndNo (netflixandno):

C I think, both forward and reverse rnxs are going on

OpenStudy (ilovebmth1234):

xD your the best okay last one lol What do we know about the consumption of an enzyme during a metabolic reaction in the body? A) The enzyme becomes part of the reactants. B) The enzyme becomes part of the products. C) The enzyme becomes part of the activated complex. D) The enzyme is not consumed and returns to its original state.

OpenStudy (n8kd_da_beast):

I think it's D. since the enzyme is only a catalyst.

OpenStudy (ilovebmth1234):

someone give him a medal??

NetflixAndNo (netflixandno):

Yeah it's D.

OpenStudy (ilovebmth1234):

thanks so much guys i appreciate it soooo much xD <3

NetflixAndNo (netflixandno):

No problem! :D

OpenStudy (n8kd_da_beast):

どういたしまして (Dōitashimapellete)you're welcome!!

OpenStudy (n8kd_da_beast):

Anytime!! :D